What is the ratio of average velocity of hydrogen molecules to that of neon atoms at the same temperature and pressure?
1
Expert's answer
2017-03-09T09:36:06-0500
Ua/Ub= SQRT(Mb/Ma)
where U is the speed of gas A and gas B;
M is molar mass of these 2 gases.
The variables for each gas are reversed on each side.
Ua/Ub=SQRT(20/2)= 3.16.

Therefore, hydrogen is about 3 times faster than neon at the same temperature and pressure.